OpenStudy (anonymous):

Integrate "square root of x times sec^2(1+x^3/2)dx? Please help or give hints?

OpenStudy (anonymous):

Try \[ y=1+x^{3/2}\quad dy=\frac 32 \sqrt x\;dx \]

OpenStudy (anonymous):

Also remember \(d(\tan x)=(\sec^2 x)\;dx\)

OpenStudy (anonymous):

Okay, so the way our professor taught us, we need to use the substitution method. So if I find u= 1+ x^3/2 and then du=3/2 sqrt x...I'm kind of confused what to do next? :/

OpenStudy (anonymous):

Do I just start taking antiderivative of sqrtxsec^2(1+x^3/2)?

OpenStudy (anonymous):

well

OpenStudy (anonymous):

\[ \sec^2(1+x^3/2) = \sec(u) \]

OpenStudy (anonymous):

Start with that.

OpenStudy (anonymous):

Okay, I remember sec(u)=sec(u)tan(u)

OpenStudy (anonymous):

\[ \sqrt x \;dx= \frac 23 \left(\frac 32 \sqrt x\;dx\right) = \frac 23 du \]

OpenStudy (anonymous):

=\[\int\limits \sqrt(x)\sec^2(1+x ^{3/2})dx\] =\[u=1+x^{3/2}\] =\[du=3/2\sqrt{x} dx \] So, =\[\int\limits 2/3(3/2\sqrt(x)dx) \sec^2(1+x ^{3/2})\] =\[\int\limits 2/3 du \sec^2(u)\] =\[=2/3\tan(1+x^{3/2}) + c\]

OpenStudy (anonymous):

okay that's what I have so far....And I know it's terribly wrong... :/ Confused..

OpenStudy (anonymous):

Question - where is the 2/3 coming from in the 4th step @aimee19

OpenStudy (anonymous):

oh, okay so this is what I was thinking. so you know how I got \[du=3/2\sqrt{x}dx\] So then I have to find a way to substitute du into the original equation, right? but there's no \[3/2\sqrt{x}dx \] in the original equation. However

OpenStudy (anonymous):

\[\sqrt{x}=\frac{ 2 }{ 3 }(\frac{ 3 }{ 2 }\sqrt{x})\]

OpenStudy (anonymous):

so I can rewrite the original equation which was \[\int\limits \sqrt(x)\sec^2(1 + x^\frac{ 3 }{ 2 }) dx \] = \[\int\limits \frac{ 2 }{ 3 }(\frac{ 3 }{ 2 }\sqrt(x) )\sec^2( 1 + x^\frac{ 3 }{ 2 })\]

OpenStudy (anonymous):

so then I can now substitute du in.

OpenStudy (anonymous):

...but I'm not sure I'm even doing the right thing. I was trying to do the substitute method to integrate the problem.

OpenStudy (anonymous):

Yeah you only use sub. when they give you a specific range to integrate on

OpenStudy (anonymous):

and the 3/2 and 2/3 cancel each other out

OpenStudy (anonymous):

you cant forget to leave out the dx, which prevents that.

OpenStudy (anonymous):

so the substitute method shouldn't be used to solve this?

OpenStudy (alekos):

yes you can still use substitution but you're just left with a constant "c" at the end of the final result. this is the best way to proceed, keep going until you have the right answer

OpenStudy (anonymous):

I'm not sure if this is right, but if I continue this is what I get: \[\int\limits 2/3 du sex^2(u) \] =\[2/3 \tan(u) + c\] =\[=2/3 \tan (1 + x^\frac{ 3 }{ 2 }) + c \]

OpenStudy (anonymous):

is that correct?

OpenStudy (alekos):

standby

OpenStudy (anonymous):

alright (:

OpenStudy (anonymous):

I'm pretty sure I messed up somewhere...

OpenStudy (alekos):

yes, i think so. i'm still working on it

OpenStudy (alekos):

i dont think it can be integrated, but i'm still looking at it

OpenStudy (alekos):

my mistake! i left something out give me another 5 minutes

OpenStudy (anonymous):

no rush! I'm basically spending the whole night working on as much review problems as I can (: I appreciate you taking the time to help!

OpenStudy (alekos):

your method was correct and the answer is definitely 2/3tan(1 + x^3/2) + c

OpenStudy (alekos):

well done!

OpenStudy (anonymous):

wait...I was actually right??!

OpenStudy (anonymous):

that's like a first for me lol >.< Thank you so much!!

OpenStudy (alekos):

if you differentiate this result you'll get the original integral!

OpenStudy (anonymous):

ah, I didn't think of checking that way. I'll use that in the future to check my answers. thanks for your help! (:

OpenStudy (alekos):

no problem you're welcome